solve ea. proportion
m-1/m-2=m-3/m+4

Answer by kiru_khandelwal(79)

(m-1)/(m-2) = (m-3)/(m+4)
=> (m-1)(m+4) = (m-3)(m-2)
=> m(m+4) -1(m+4) = m(m-2) -3(m-2)
=> m^2 + 4m -m -4 = m^2 -2m -3m + 6
=> m^2 + 3m -4 = m^2 -5m + 6
=> 3m -4 = -5m + 6
=> 3m + 5m = 6 + 4
=> 8m = 10
=> m = 10/8
=> m = 5/4